# SandCell: Sandboxing Rust Beyond Unsafe Code

### v1

@@ 7.3 In-Process Isolation
. To ensure that heap-allocated objects are placed into memory belonging to a particular domain, SDRaD provides a domain-aware memory allocator built on the Two-Level Segregated Fit (TLSF) heap allocator [51]. However, TLSF implementation used in SDRaD only supports 8-byte alignment for allocations, whereas certain common Rust crates, such as hashbrown, used by the cargo test and cargo bench tools expect allocations to be made at 16-byte alignment. SDRaD-v2 replaces TLSF with the mimalloc general-purpose allocator [45]. In addition, the TLSF-based SDRaD allocator places allocator metadata on the domain heap, where it could be corrupted by memory errors that occur within sandboxed code. The mimalloc allocator, similarly, places allocator metadata in a “backheap" area, that by default would be placed on the domain heap. To harden the SDRaD-v2 allocator, we modified mimalloc to ensure that all metadata is allocated exclusively from the monitor domain, rather than from a domain’s backheap. This guarantees that allocator metadata is restricted to the monitor domain, preventing any modifications from sandboxed code. In SDRaD-v2, the allocator algorithm that is run when malloc or sdrad_malloc is called (➅ in Figure˜3), always runs in the monitor domain (➃).
@@
7.4 In-process System Call Filtering.
SandCell needs a solution for system call interposition 1) to harden PKU-based in-process isolation and 2) to reduce the attack surface for each SandCell domain. Early research on hardening PKU-based isolation using system-call interposition can either suffer from high overhead or offer kernel changes or kernel modules [60, 61, 57]. To address these limitations, we revisited existing system call interposition methods and employed a newer approach, zpoline [65]: a technique for the x86-64 architecture to intercept system calls. It replaces the two-byte syscall or sysenter instructions with a call rax instruction of equal length. Since the system call number is stored in the rax register as per the x86-64 call conventions, the call rax causes a control-flow transfer to a memory address corresponding to the system call number, within the first 500 bytes of memory. zpoline populates this range with a series of no-operation instructions, effectively constructing a NOP sled, which eventually leads to a trampoline transferring control to a user-defined code or a syscall interposer’s code. This design ensures that the rewriting of system call instructions cannot fail, providing a reliable method for intercepting and customizing system call behavior in applications. zpoline handles static binaries before main, but it doesn’t support dynamically generated code. To address this, we hook the mprotect call that marks memory as executable and scan the memory region for syscall instructions and rewrite them. This allows zpoline to trace syscalls in dynamically generated code as well. Our system call policy enforcement, implemented on top of zpoline, consists of approximately 300 lines of code C/C++ code in total.
@@
SDRaD-v2 enforces similar system call filter policies as previous work on PKU-Based isolation primitives [60, 61]. Using the system call interpositioning primitive, zpoline, (➇ in Figure˜3) we trap system calls from any sandbox domains to a system call filtering engine in the monitor domain (➆). The policy engine applies a default filtering policy that rejects any system calls that attempt to circumvent the SDRaD-enforced PKU policy, such as pkey_mprotect calls that attempt to change the PKU keys associated with memory pages, or open family systems calls to /proc/self/mem which would give a sandbox raw access to the memory of the process. SDRaD-v2 also enforces that pages being dynamically mapped as executable do not contain xrstor and wrpkru instructions which could be used to bypass PKU policies or alter the monitor state. To ensure this, the monitor scans such pages for the presence of these instructions, and enforces a Write-XOR-Execute policy on executable pages after inspection. Also, any system call that attempts to modify the zpoline NOP sled area is rejected by our enforcement mechanism. Systems calls from the monitor domain itself are exempt from this policy to enable to SDRaD domain management logic (➄) to manage the sandbox PKU policies properly.
7.5 TCB Size
@@
At runtime, the trusted computing base includes the memory allocator, the SDRaD monitor, and the generated context-switch code. The memory allocator contains $\approx$10k lines of code. The SDRaD monitor contains $\approx$5k lines of code. The generated context-switch code contains around 20 to 100 lines of code for each sandboxed function call.
8 Evaluation
In this section, we evaluate SandCell on several Rust applications. First, among the Rust crates that have had memory safety vulnerabilities recorded in the RustSec Advisory Database and whose interfaces do not use the currently unsupported language features described in §˜10, we randomly selected four crates for evaluation, including transpose, rouille, elf-rs and rust-openssl. To evaluate SandCell on a larger application, we also applied it to Servo, an experimental web browser engine in Rust, and ripgrep, a replacement for grep.
For each application, we demonstrate how we sandboxed the application, the required programmer effort, runtime overhead introduced by sandboxing, and how our sandboxcing can prevent real-world vulnerabilities. For all applications, if applicable, we compare the difference between copy-only and heap-sharing to show the effectiveness of SandCell’s heap-sharing optimization. We also compare the performance of SandCell with SandCrust [42] and SDRaD-FFI [34].
@@ 8 Evaluation
We ran all experiments on a 48-core Intel Xeon Gold 5318S CPU running at 2.10 GHz and 196 GiB of RAM. The machine runs Ubuntu 22.04.3 LTS with version 5.15.0-83 of the Linux kernel. We use rustc version nightly-2024-02-03.
tranpose rouille elf-rs ripgrep rust-openssl Servo
Spec Lines 2 2 2 2 2 2
Table 1: Lines of SandCell specification for applications.
Sandboxed SandCell-Copy SandCell-Share Boundary
LOC Max Min Max Min Size
transpose 671 3.84% 1.91% 0.47% -0.64% 6
rouille 10,422 26.43% -1.40% 17.49% -2.75% 1
elf-rs 1,342 25.63% 1.65% 3.39% 0.78% 1
ripgrep 2,736 - - 28.4% 18.2% 8
rust-openssl 49,306 13.12% 2.62% 3.38% -0.71% 28
Servo 134,708 - - 10.06% -0.86% 1
@@
Table 2: SandCell Performance maximum and minimum overhead compared to baseline. Missing data indicates the interface contains unsupported Rust features (see §˜10). Max and min are maximal and minimal overheads under different benchmarking workloads, respectively; see tables in the appendix for details.
8.1 User Effort for Specification
To write specification for SandCell, we need to identify the isolation boundaries in the application. In the evaluation, we explore two strategies of finding isolation boundaries: library-based and user-based. For library-based isolation, we look for crates with known vulnerabilities or known to be error-prone, and sandbox them entirely. We apply this strategy to transpose, ripgrep, elf-rs and rust-openssl. Writing a specification using the library-based strategy requires little user effort; only knowing which library to isolate is sufficient.
For user-based isolation, we isolate code shared by different users, mitigating potential attacks from one user to another. We have rouille and Servo under this category. In this case, we need to follow the code path starting from the user input to the component that processes the input, then mark that component as isolated. This requires more effort compared to the library-based approach, but is still manageable. For example, in the rouille web server, we traced the code path from the HTTP request handler to the compression function calling zlib, which involves 3 function calls in between.
After determining the isolation boundaries, writing the corresponding specification is straightforward. We have the specification size for each benchmark in Table˜1. Notice that each application requires only two lines of specification.
8.2 Applications
In this subsection, we introduce all the applications used in the evaluation. For each application, we describe known vulnerabilities, the sandboxing architectures that can mitigate these vulnerabilities, our evaluation methods, and performance results. In the main text we summarize statistics of applications in Table˜2, including the lines of code in sandboxes, summary performance results, the lines of code in the policy specification, and the number of functions exposed by sandboxes (boundary size). The detailed performance data can be found in the appendix from Table˜3 to Table˜8.
Figure 6: Performance overhead of SandCell on Servo for different categories in Dromaeo benchmark.
Servo
. As a browser engine, Servo [31] needs to spawn a JavaScript execution thread for each browser tab to run potentially dangerous JavaScript code. To mitigate the risk posed by malicious JavaScript code, Servo provides a multiprocess mode, isolating all JavaScript execution threads in a separate process. However, cross-process data transmission introduces performance overhead. Furthermore, there is no isolation between individual JavaScript threads within the isolated process. In our experiments, we created a separate sandbox for each JavaScript execution thread. This design keeps all threads within the main process while maintaining in-process memory isolation that prevents a malicious script in one JavaScript thread from attacking other threads. Threads in Servo use channels to communicate, which involve heap allocations. We modified the channel implementation by adding 300 lines of code to make it domain-aware, so that SandCell can change their allocation sites to shared domains and allow communication between sandboxes.
We measured the performance overhead of SandCell on Servo with the Dromaeo benchmark included in Servo’s official repository. The benchmark includes a set of JavaScript tests that are designed to measure the performance of various JavaScript operations. For each test, the benchmark runs it for 5 times and takes the average of the running time. Tests are grouped into 5 categories and the average overhead for each category is shown in Figure˜6. Overall, the average overhead is 3.1%.
@@ 8.2 Applications
OpenSSL [7] is a widely-used cryptographic library. In Rust, a crate rust-openssl provides a wrapper around OpenSSL and is widely used in networking-related Rust applications. OpenSSL has a history of vulnerabilities, such as CVE-2023-0215, CVE-2023-0216, and CVE-2023-0217. Moreover, rust-openssl itself introduces vulnerabilities. For example, RUSTSEC-2023-0044 is a vulnerability in which rust-openssl passes an empty string to OpenSSL functions without properly setting null byte termination. Then it causes an out-of-bound read when one of the OpenSSL functions calls strlen on the string.
In our evaluation, the interface between the HTTPS client and rust-openssl is relatively large with 28 functions and it is not easy to sandbox by hand using tools like SandCrust or SDRaD-FFI. We use SandCell to sandbox the entire rust-openssl crate.
We measure the overhead by running an HTTPS client that uses rust-openssl to do a GET request and a POST request to a server. We measure the running time of the client for both requests and compare the performance of SandCell with the baseline. The results show that SandCell-Copy introduces a maximum overhead of 18.56% and a minimum of -0.35%, while SandCell-Share introduces a maximum overhead of 3.38% and a minimum of -0.71%.
@@
The transpose [11] crate provides functions to transpose matrices represented as vectors. Users are required to pass the height and width of the input matrix to these functions, which then check that the vector’s length equals the product of the height and width. However, in vulnerability RUSTSEC-2023-0080 [14], the product of the height and width could overflow the integer type. In debug mode, the program aborts upon overflow; in release mode, however, the product wraps around, allowing the program to continue running. This results in an out-of-bounds write when accessing the vector based on the incorrect height and width. For benchmarking, we use the RustFFT crate, which relies on transpose and implements various Fast Fourier Transform (FFT) algorithms. We sandbox the transpose function to prevent any potential exploitation of the out-of-bounds write from affecting the rest of the program. Then, we run FFT algorithms on matrices of varying sizes to measure the performance overhead introduced by sandboxing. We used transpose version 0.2.3 for our evaluation. Our results show that SandCell can introduce a maximum overhead of 3.84% when data is copied between domains over the unsandboxed baseline. Nevertheless, when data is shared between domains, it results in a maximum overhead of 0.47% and a minimum of -0.64%. The negative overhead is because the data transmission overhead is dwarfed by the FFT computation, when the matrix size is large.
@@
rouille [6] is a web server framework that provides a simple API to handle HTTP requests. To handle HTTP compression, we modified rouille to use the flate2 crate, which provides a wrapper around the zlib library. zlib is written in C and known to have vulnerabilities, such as CVE-2018-25032 [26], CVE-2022-37434 [28] and CVE-2002-0059 [25]. We implement a web server using rouille that sends a fixed string as response to all requests. Then we sandbox the compress function called by rouille to prevent the vulnerabilities in zlib from being exploited. We measure the performance degradation in rouille by comparing latency and throughput of the web server over different thread configurations. We used the wrk client with 10 threads, with the same static resource for 10 seconds and repeating the test 10 times. SandCell-Copy introduces a maximum overhead of 26.43% and a minimum of -1.40%, while SandCell-Share introduces a maximum overhead of 17.49% and a minimum of -2.75%. The performance overhead is not positively correlated with the number of concurrent connections, but in general SandCell-Copy performs worse because of the overhead of copying data to the sandboxed compression library.
@@
elf-rs is a crate that provides functions to parse and manipulate ELF files. In an ELF file, an offset field is stored at a fixed place in the file header. The ELF parser is required to follow the offset to read the actual content of the file. However, elf-rs vulnerability RUSTSEC-2022-0079 [13] failed to check the offset was within the file size, leading to an out-of-bound access. We implement a readelf tool using elf-rs and sandbox the parsing function from elf-rs to prevent the out-of-bound access from touching other parts of the program. The running time of readelf is measured by parsing different ELF files with different sizes. We measured the parsing time of each file, comparing the performance of SandCell with the unsandboxed baseline over $10^{5}$ iterations. SandCell-Copy introduces a maximum overhead of 25.63% and a minimum of 1.65%, while SandCell-Share introduces a maximum overhead of 3.39% and a minimum of 0.78%. The overhead of SandCell-Copy goes higher when the file size gets larger, while the overhead of SandCell-Share is relatively stable because the size of data crossing the boundary does not change much with larger files.
@@
ripgrep [2] is a replacement for grep written in Rust, which is a Linux utility for regular-expression pattern matching in files. An important component of ripgrep is the glob pattern matching for file names, which is prone to errors and directly interacts with untrusted data. We isolate this component by specifying a sandbox isolating the GlobSet type, which is responsible for the glob pattern matching. Our copying implementation cannot handle this benchmark as discussed in §˜10; so we evaluate the performance of only the SandCell-Share implementation. We measure the running time of ripgrep by searching for a fixed pattern for different numbers of files. Compared with the baseline, SandCell-Share introduces a maximum overhead of 13.12% and a minimum of 2.62%.
Figure 7: Execution Time of compress() and uncompress() for SandCell, SandCrust, SDRaD-FFI with SDRaD-v1, SDRaD-FFI with SDRaD-v2, Baseline, and Baseline with mimalloc allocator
Comparison with related work.
We compared SandCell with SDRaD-FFI and SandCrust using the compression library snappy and the image codec libpng. For fairness, we reused the public benchmark artifact from SDRaD-FFI, which measures compression and decompression times on randomly generated data of varying sizes over $5\times 10^{5}$ iterations. Figure˜7 presents the overall execution times across different data sizes. As a baseline, we evaluated both the standard libc allocator and mimalloc. While the difference is negligible for small buffer sizes, mimalloc provides a clear advantage for larger inputs. Since SandCell uses mimalloc by default, this explains why SandCell outperforms the libc baseline for larger inputs. We also measured SDRaD-FFI with its default TLSF allocator and our SDRaD-v2 variant, which enhances compatibility and security. As expected, SDRaD-v2 incurs additional overhead compared to SDRaD-FFI, primarily due to the allocator choice and the syscall interposition mechanism. SandCrust, which enforces process-based isolation, shows slower performance than SandCell because of the overhead of process context switching and inter-process communication. Finally, although SDRaD-FFI provides in-process isolation, it still performs worse than SandCell because its data transfers rely on explicit copying across domains, whereas SandCell reduces this overhead through shared heap access.
Discussion.
By comparing the SandCell-Copy with SandCell-Share and comparing SandCell with SDRaD-FFI, we show that the heap-sharing optimization can effectively reduce the overhead of sandboxing. Also, by comparing with SandCrust, we show in-process isolation can provide better performance than process-based isolation. Table˜2 also lists the user effort required when using SandCell in the form of numbers of specification lines. The specification just needs to list the relevant crate names and syscalls; other than rust-openssl, none of other user cases require syscalls in sandboxes. This shows that SandCell can provide a lightweight sandboxing mechanism with which users can achieve desired security with low effort.
9 Related Work

### v2

@@ 7.3. In-Process Isolation
. To ensure that heap-allocated objects are placed into memory belonging to a particular domain, SDRaD provides a domain-aware memory allocator built on the Two-Level Segregated Fit (TLSF) heap allocator (Masmano04, ). However, TLSF implementation used in SDRaD only supports 8-byte alignment for allocations, whereas certain common Rust crates, such as hashbrown, used by the cargo test and cargo bench tools expect allocations to be made at 16-byte alignment. SDRaD-v2 replaces TLSF with the mimalloc general-purpose allocator (Leijen2019MimallocFL, ). In addition, the TLSF-based SDRaD allocator places allocator metadata on the domain heap, where it could be corrupted by memory errors that occur within sandboxed code. The mimalloc allocator, similarly, places allocator metadata in a “backheap” area, that by default would be placed on the domain heap. To harden the SDRaD-v2 allocator, we modified mimalloc to ensure that all metadata is allocated exclusively from the monitor domain, rather than from a domain’s backheap. This guarantees that allocator metadata is restricted to the monitor domain, preventing any modifications from sandboxed code. In SDRaD-v2, the allocator algorithm that is run when malloc or sdrad_malloc is called (➅ in Figure 3), always runs in the monitor domain (➃).
@@
7.4. In-process System Call Filtering.
SandCell needs a solution for system call interposition 1) to harden PKU-based in-process isolation and 2) to reduce the attack surface for each SandCell domain. Early research on hardening PKU-based isolation using system-call interposition can either suffer from high overhead or offer kernel changes or kernel modules (Vahldiek-Oberwagner19, ; Voulimeneas22, ; Schrammel20, ). To address these limitations, we revisited existing system call interposition methods and employed a newer approach, zpoline (yasukata23, ): a technique for the x86-64 architecture to intercept system calls. It replaces the two-byte syscall or sysenter instructions with a call rax instruction of equal length. Since the system call number is stored in the rax register as per the x86-64 call conventions, the call rax causes a control-flow transfer to a memory address corresponding to the system call number, within the first 500 bytes of memory. zpoline populates this range with a series of no-operation instructions, effectively constructing a NOP sled, which eventually leads to a trampoline transferring control to a user-defined code or a syscall interposer’s code. This design ensures that the rewriting of system call instructions cannot fail, providing a reliable method for intercepting and customizing system call behavior in applications. zpoline handles static binaries before main, but it doesn’t support dynamically generated code. To address this, we hook the mprotect call that marks memory as executable and scan the memory region for syscall instructions and rewrite them. This allows zpoline to trace syscalls in dynamically generated code as well. Our system call policy enforcement, implemented on top of zpoline, consists of approximately 300 lines of code C/C++ code in total.
@@
SDRaD-v2 enforces similar system call filter policies as previous work on PKU-Based isolation primitives (Vahldiek-Oberwagner19, ; Voulimeneas22, ). Using the system call interpositioning primitive, zpoline, (➇ in Figure 3) we trap system calls from any sandbox domains to a system call filtering engine in the monitor domain (➆). The policy engine applies a default filtering policy that rejects any system calls that attempt to circumvent the SDRaD-enforced PKU policy, such as pkey_mprotect calls that attempt to change the PKU keys associated with memory pages, or open family systems calls to /proc/self/mem which would give a sandbox raw access to the memory of the process. SDRaD-v2 also enforces that pages being dynamically mapped as executable do not contain xrstor and wrpkru instructions which could be used to bypass PKU policies or alter the monitor state. To ensure this, the monitor scans such pages for the presence of these instructions, and enforces a Write-XOR-Execute policy on executable pages after inspection. Also, any system call that attempts to modify the zpoline NOP sled area is rejected by our enforcement mechanism. Systems calls from the monitor domain itself are exempt from this policy to enable to SDRaD domain management logic (➄) to manage the sandbox PKU policies properly.
7.5. TCB Size
@@
At runtime, the trusted computing base includes the memory allocator, the SDRaD monitor, and the generated context-switch code. The memory allocator contains $\approx$10k lines of code. The SDRaD monitor contains $\approx$5k lines of code. The generated context-switch code contains around 20 to 100 lines of code for each sandboxed function.
8. Evaluation
We evaluate SandCell along three dimensions: security, usability, and performance. For security, we demonstrate that SandCell mitigates real-world vulnerabilities. For usability, we apply SandCell to real Rust applications and show that isolation policies can be expressed with little specification effort. For performance, the dominant cost is sandbox context switching; so overhead depends on program behavior and boundary placement: frequent crossings amplify overhead, while boundaries on infrequent paths (e.g., one-shot parsing) incur little overhead. To understand how overhead varies across usage patterns, we first measure per-call switching cost in micro-benchmarks against prior tools and the baseline. With that cost in hand, we apply SandCell to diverse real-world applications with different call patterns to quantify end-to-end overhead. Specifically, we answer the following research questions:
  • RQ1 (User Effort): How much effort is required to write specifications for SandCell?
  • RQ2 (Security): Can SandCell effectively mitigate memory safety vulnerabilities?
  • RQ3 (Micro-benchmarks): How does SandCell compare with existing sandboxing tools in controlled benchmarks?
  • RQ4 (Real-world Overhead): What is the runtime overhead when applying SandCell to real-world applications?
Benchmarks.
Across the evaluation, we use two classes of real-world Rust programs: crates with known memory-safety advisories and larger applications/parsing libraries. We randomly selected four Rust crates with memory safety vulnerabilities recorded in the RustSec Advisory Database: transpose (transpose, ), rouille (rouille, ), elf-rs (elfrs, ), and rust-openssl (openssl, ). We also applied SandCell to larger applications and parsing-heavy libraries, including Servo (servo, ), ripgrep (githubGitHubBurntSushiripgrep, ), oq (oq, ), and async-graphql (asyncgraphql, ).
Experiment setup.
@@ 8. Evaluation
We ran all experiments on a 12-core 13th Gen Intel(R) Core(TM) i7-1360P at 2.10 GHz and 64 GiB of RAM. The machine runs Ubuntu 24.04.3 LTS with version 6.12.25 of the Linux kernel. We disable hyperthreading on the CPU to reduce measurement noise.
8.1. RQ1: User Effort for Specification
Application Spec Lines Sandboxed LOC Boundary
transpose 2 671 6
elf-rs 2 1,342 1
rust-openssl 2 49,306 28
rouille 2 10,422 1
Servo 2 134,708 1
ripgrep 2 2,736 8
oq 2 1,767 5
async-graphql 2 2,465 1
Table 1. Specification size, sandboxed LOC, and public API numbers for benchmark applications or libraries. Notice that here we only count the lines of code for the target crate itself, excluding its dependencies. Although at runtime, the sandboxed code may include code from dependencies as well.
We apply the specification strategies described in § 5 to our benchmark applications. Table 1 summarizes the results.
Library-based isolation.
We apply this strategy to three crates with known memory-safety vulnerabilities:
  • transpose is a matrix library with an out-of-bounds write (RUSTSEC-2023-0080 (rustsecRUSTSEC20230080Transpose, )) triggered by crafted input dimensions.
  • elf-rs is an ELF file parser with a heap buffer overflow (RUSTSEC-2022-0079 (RUSTSEC-2022-0079, )) when parsing malformed binaries.
  • rust-openssl wraps OpenSSL, which has had numerous vulnerabilities including CVE-2023-0216 (CVE-2023-0216, ) (invalid pointer dereference) and RUSTSEC-2023-0044 (rustsecRUSTSEC20230044Openssl, ) (use-after-free).
User-based isolation.
We apply this strategy to two multi-tenant applications:
  • rouille is an HTTP server framework. It uses zlib for compression, which had vulnerabilities including CVE-2018-25032 (CVE-2018-25032, ), CVE-2022-37434 (CVE-2022-37434, ), and CVE-2002-0059 (CVE-2002-0059, ). We isolate each call to the compression function for each user, preventing a compromised request from accessing other users’ data. Identifying this boundary required tracing through 3 functions from the HTTP request handler.
  • Servo is Mozilla’s experimental browser engine. We isolate the JavaScript execution engine for each tab, preventing a malicious script in one tab from accessing data in other tabs. We identified the script execution component and its entry function from the project’s architecture documentation.
While this strategy requires understanding the application’s data flow, the effort remains manageable.
Input-oriented isolation.
We apply this strategy to three applications that process external input:
  • ripgrep is a fast grep tool. We isolate its GlobSet struct, which matches file names against user-provided glob patterns. We identified this by examining the dependency tree, finding the globset crate, and locating its public interface struct.
  • oq is a command-line tool for querying and transforming data in various formats (JSON, YAML, TOML). We isolate the entire oq crate, which parses untrusted input files.
  • async-graphql is a GraphQL server library. We isolate its parsing function, parse_query, which parses untrusted client queries. The identification was straightforward due to the clear function name.
These components directly handle external input and benefit from isolation regardless of whether there are currently known vulnerabilities.
Specification effort.
As shown in Table 1, each application requires only two lines of specification. Once the isolation boundary is identified—whether a crate, struct, or function—specifying it in SandCell is straightforward. The main effort lies in understanding the application’s structure to identify appropriate boundaries, not in writing the specification itself.
8.2. RQ2: Security Evaluation
We evaluate SandCell’s security effectiveness through two complementary approaches: (1) analyzing how SandCell would contain known CVEs in real-world crates, and (2) validating isolation guarantees using proof-of-concept (PoC) exploits derived from cross-language attack patterns.
Analysis of known vulnerabilities.
Our benchmark crates contain known vulnerabilities including RUSTSEC-2023-0080 (rustsecRUSTSEC20230080Transpose, ), CVE-2018-25032 (CVE-2018-25032, ), RUSTSEC-2023-0044 (rustsecRUSTSEC20230044Openssl, ), and RUSTSEC-2022-0079 (RUSTSEC-2022-0079, ) (out-of-bound writes), as well as RUSTSEC-2023-0044 (rustsecRUSTSEC20230044Openssl, ), CVE-2002-0059 (CVE-2002-0059, ) and CVE-2023-0216 (CVE-2023-0216, ) (double-free and invalid pointer dereference). By sandboxing the vulnerable function or crate with SandCell, these vulnerabilities would be contained: if triggered, exploits can only access memory inside the sandbox or will be caught by the low-level isolation mechanism of SandCell, which throws a runtime error when memory outside the sandbox is accessed.
PoC-based validation.
To concretely validate SandCell’s isolation guarantees, we developed three proof-of-concept exploits modeling attack patterns from prior work (claPaper, ; mixedBinariesPaper, ), including both pure-Rust and cross-language attacks:
  • Lifetime soundness exploit: A pure-Rust attack exploiting the lifetime expansion soundness hole (Rust issue #25860 (rust25860, )) to forge a mutable reference to an arbitrary address. This models supply-chain attacks where safe Rust can corrupt memory due to compiler unsoundness. When run inside SandCell’s sandbox, cross-domain writes to the monitor cause memory faults handled by SDRaD-v2, leaving the target unchanged.
  • Lifetime violation via C FFI (claPaper, ): Rust allocates a buffer, hands its pointer to C, and C frees it behind Rust’s back; Rust then reads the freed memory, creating a use-after-free. Under SandCell, when the sandbox attempts to free a buffer allocated by the monitor domain, the monitor intercepts and blocks the operation, preventing the attacker from invalidating host memory.
  • Mixed stack corruption (mixedBinariesPaper, ): When Rust and C coexist, Rust stack frames are not protected by SafeStack; a buffer overflow in C can write past a Rust stack buffer and corrupt adjacent control data (e.g., a return address/guard), enabling control-flow hijack. After isolation, the untrusted C runs inside a sandbox domain; so the overflow corrupts only the sandbox’s own stack; the monitor’s control data remains intact, containing the attack’s impact.
Each PoC includes an attack binary demonstrating the exploit without isolation, and an isolated binary showing mitigation under SandCell. In all cases, cross-domain memory accesses from sandbox code into monitor memory result in runtime faults, successfully containing the attack.
8.3. RQ3: Micro-benchmark Comparison
Figure 5. Execution Time of compress() and uncompress() for SandCell, SandCrust, SDRaD-FFI with SDRaD-v1, SDRaD-FFI with SDRaD-v2, Baseline, and Baseline with mimalloc allocator
We measure per-call overhead using the compression library snappy and the image codec libpng, comparing SandCell against SDRaD-FFI (gulmezFriendFoeExploring2023, ) and Sandcrust (lamowskiSandcrustAutomaticSandboxing2017, ).
Benchmark construction.
Each benchmark program executes exactly one sandboxed function call per iteration, measuring total execution time (sandbox overhead plus workload processing). For snappy, we wrap the compress and uncompress functions, invoking each once per iteration while varying the input buffer size (1B to 2MB, in powers of two). For libpng, we wrap decode_png and invoke it once per iteration on different test images. By varying input size, we observe how fixed sandbox overhead amortizes as workload increases. We perform 100 warm-up iterations followed by 50,000 timed iterations per configuration, measuring wall-clock time and computing mean and standard deviation. We reused the public benchmark artifact from SDRaD-FFI (gulmezFriendFoeExploring2023, ) to ensure a fair comparison. Figure 5 presents the overall execution times across different data sizes.
Baseline comparison.
We evaluated both the standard libc allocator and mimalloc as baselines. While the difference is negligible for small buffer sizes, mimalloc provides a clear advantage for larger inputs. Since SandCell uses mimalloc by default, this explains why SandCell outperforms the libc baseline for larger inputs.
SDRaD-FFI comparison.
We measured SDRaD-FFI with its default TLSF allocator and our SDRaD-v2 variant, which enhances compatibility and security. As expected, SDRaD-v2 incurs additional overhead compared to SDRaD-FFI, primarily due to the allocator choice and the syscall interposition mechanism. Although SDRaD-FFI uses in-process isolation, it performs worse than SandCell because its data transfers rely on copying across domains, whereas SandCell reduces this overhead through shared heap access.
Sandcrust comparison.
Sandcrust, which uses processes for isolation, shows worse performance than SandCell due to the overhead of process context switching and inter-process communication.
Summary.
These micro-benchmarks demonstrate that: (1) the heap-sharing optimization of SandCell effectively reduces sandboxing overhead compared to copy-based approaches (SDRaD-FFI); and (2) in-process isolation performs better than process-based isolation (Sandcrust).
8.4. RQ4: Real-world Application Overhead
SandCell-Copy SandCell-Share
Min Max Min Max
rouille 37.1% 89.0% -0.3% 3.0%
elf-rs 12.7% 96.9% 5.4% 15.8%
oq 5.7% 36.0% 8.4% 33.2%
async-graphql 6.8% 15.8% 5.3% 13.3%
ripgrep 5% 13.4% 2.9% 10.3%
rust-openssl 3.9% 17.7% 0.4% 22.2%
Servo - - -1.4% 9.4%
transpose 3.3% 4.4% 3.0% 4.8%
@@ 8.4. RQ4: Real-world Application Overhead
Table 2. SandCell performance overhead compared to the baseline. Missing data indicates the application requires data sharing and cannot be accommodated by the copy mode. Max and min are maximal and minimal overheads under different workloads. For rouille, we use the throughput here. See Appendix B for details.
We evaluate SandCell’s overhead on eight real-world applications and libraries. We compare SandCell-Copy (data copied between domains) and SandCell-Share (heap-sharing optimization). Table 2 summarizes the performance overhead ranges. Detailed performance data can be found in the appendixAppendix B.
Servo.
We run the Dromaeo JavaScript benchmark suite, which executes thousands of JavaScript operations per test category. Each operation triggers a sandbox entry/exit as the JS engine runs in isolation. The high call frequency (hundreds of calls per millisecond) makes this a stress test for sandbox transition cost. Threads in Servo use channels to communicate, which involve heap allocations. We modified the channel implementation by adding 300 lines of code to make it domain-aware, so that SandCell can change their allocation sites to shared domains and allow communication between sandboxes. Copy mode is not applicable because the inter-thread message channel requires real memory sharing. Overall, the max overhead is 9.4%.
@@
We use an HTTP library in Rust, attohttpc (attohttpc, ) to implement seven request-level tasks (get, post, session, assets, multipart, redirect, stream). Together they cover basic GET and POST, a session with multiple requests and JSON parsing, asset-style transfers with HEAD and ETag conditional GET, multipart upload of a  256 KB payload, redirect chains, and a streamed JSON response; all responses are consumed and checksummed to keep work on the hot path. These tasks are executed against public endpoints (e.g., www.google.com and httpbin.org) to exercise realistic HTTP client behavior. These program can be found in our artifacts (appendix A). Each TLS handshake and data transfer invokes multiple sandboxed OpenSSL functions. SandCell-Copy introduces 17.7.0% max overhead; SandCell-Share reduces this to 22.2% max.
@@
We integrate the sandboxed transpose function into RustFFT, which uses matrix transpositions as part of FFT computation. We vary matrix sizes from $\times64$ to $\times4096$. Each FFT operation triggers one sandbox call, and data transfer for SandCell-Copy scales quadratically with matrix size. For performance measurement, we run the FFT algorithm benchmark that comes with RustFFT. Both SandCell-Copy and SandCell-Share have 3% to 5% overhead because the FFT computation dominates the running time.
@@
We run a local rouille server with a configurable thread pool and a single handler that returns a fixed‑size response body, and the response passes through rouille’s content encoding where compression work is done by the sandboxed function, which calls zlib. We then drive that server with wrk. We configure wrk with 8 threads and 200 connections. During the experiement, we run rouille with 1,2 and 4 threads. Each request triggers one sandbox call to the compress function. We measure both latency and throughput. SandCell-Copy introduces 89% max overhead for latency and 77.2% for throughput; SandCell-Share reduces this to 14.9% max and 3% max.
@@
We build a readelf-like tool that parses ELF binaries and prints its content. We test on different sizes of binaries. Each file triggers one sandbox call to the parser, with data transfer proportional to file size. We measure the running time for each file. SandCell-Copy introduces 96.9% max overhead; SandCell-Share reduces this to 15.8% max.
oq.
The oq benchmark evaluates the oq CLI (a jq-like object query tool) by running a fixed filter over JSON inputs of increasing size. For each sample file (100k–500k JSON) that contains an items array, we execute oq to parse the input and count the array length. The size of data across the boundary for SandCell-Copy is proportional to the input size. For smaller files, the parsing time is dominated by the sandbox transition overhead, therefore, both SandCell-Copy and SandCell-Share show higher overhead (36.0% and 33.2% respectively). When the input size increases, the parsing time dominates, and the overhead decreases (5.7% and 8.4% respectively).
async-graphql.
We setup a GraphQL server using async-graphql with a synthetic dataset (including 2000 entities and relations between them). Then we measure the end-to-end performance with four kinds of queries, covering fragments, pagination/edges, conditional directives, and nested relationships. Each incoming query triggers one sandbox call to parse_query. The parsed AST is returned to the monitor for execution. SandCell-Copy introduces 15.8% max overhead; SandCell-Share reduces this to 13.3% max.
@@
We use ripgrep to search directory trees with varying numbers of files (64–512). To check pattern matches, the methods of the isolated struct GlobSet are called several times per file , while the array of matched patterns is sent across the boundary. We measure the end-to-end running time. SandCell-Copy introduces 13.4% max overhead; SandCell-Share reduces this to 10.3% max. However, the rest of the overhead is not from the context switching, instead, it is because of the zpoline syscall interception mechanism has non-negligible overhead for the open syscall, which ripgrep invokes frequently when searching files.
Summary.
Overhead correlates with call frequency and data transfer volume. High-frequency call patterns (e.g., rust-openssl) stress sandbox transition costs, while large data transfers (e.g., elf-rs) stress the copying mechanism. SandCell-Share outperforms SandCell-Copy by eliminating data copying overhead in most cases. In some cases, the data size across isolation boundary is small and the noise dominates, so that SandCell-Share performs similar or slightly worse than SandCell-Copy.
9. Related Work